Equation of straight line is 2x + 3y = 5 Then how to find the slope of straight line? 

Grade:11

1 Answers

Arun
5 years ago
Dear student
Try to convert it in the form y = mx +c
Hence
2x + 3y = 5
3y = -2 x +5
y = -2x/3 + 5/3
Hence comparing with y = mx +c
m = -2/3 which is the slope